Carter Company acquired three machines for $560,000 in a package deal. An appraisal indicated that the three machines had the following market values: Machine 1: $180,000 Machine 2: $150,000 Machine 3: $360,000 What is the cost for Machine 2 for Carter Company?
Answer #1
Machine Market values Percentage of market values Allocated cost
Machine 1 $180,000 180,000/690,000 = 0.260869565 560,000 x 0.260869565 = 146,087
Machine 2 $150,000 150,000/690,000 = 0.217391304 560,000 x 0.217391304 = 121,739
Machine 3 $360,000 360,000/690,000 = 0.52173913 560,000 x 0.52173913 = 292,174
Total $690,000 $560,000

cost for Machine 2 for Carter Company = $121,739

Please ask if you have any query related to the question. Thank you